I cannot seem to solve the following question:

〖(2+2i)〗^13

Answer by Alan3354(69443)

(2+2i)^13
= (sqrt(8)cis(45))^13
= 8^6*sqrt(8)cis(585)
= 524288*sqrt(2)cis(225)
= -262144 - 262144i